OpenStudy (anonymous):

Find all the zeros of each polynomial function. f(x) = x4 – 5x2 – 36 Select one: a. ±3, ±2i b. ±3 only c. ±2i only d. none of the above

OpenStudy (anonymous):

@rational

jhonyy9 (jhonyy9):

note x^2 = t and will get a quadratic equation of t

OpenStudy (anonymous):

try to reduce it first to quadratic since the exponents are even and trinomial we have... say x=u^2... so equation is converted to f(u) = u^2 - 5u -36... can apply quadratic formula

jhonyy9 (jhonyy9):

yes right but than you see it now you can factoriz it yes ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes you can... :)

OpenStudy (anonymous):

so which would it be

jhonyy9 (jhonyy9):

f(u)=(u-9)(u+4) right ?

OpenStudy (anonymous):

correct

jhonyy9 (jhonyy9):

so do you can continue it now ?

OpenStudy (anonymous):

sure

jhonyy9 (jhonyy9):

ok was my pleasure

OpenStudy (anonymous):

i think its a actuallu

jhonyy9 (jhonyy9):

si do you can calcule it here please ?